# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 3.1 A message passing algorithm
$w(u,v)>Z_{h}(u,v)+Z_{h}(v,u)\Rightarrow\{v\}=\operatorname*{argmax}_{v^{\prime}\sim v}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})).$
Figure 5: Construction of the law of $Z_{h}$ on a $1-$neighbourhood.
@@
To show the equivalent rule, we only need to show that $u$ is not matched if and only if $\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\leq 0$. For the implication, we have that any incident $Z_{h}(v,u)$ for $v\sim u$ evaluates to zero by applying Recursion (3). So for any $v\sim u$, $Z_{h}(v,u)+Z_{h}(u,v)=Z_{h}(u,v)$. Since $\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\leq 0$, we deduce that for any $v^{\prime}\sim u$, $w(u,v^{\prime})\leq Z_{h}(u,v^{\prime})$ , but we have just seen that $Z_{h}(v^{\prime},u)=0$ hence $w(u,v^{\prime})\leq Z_{h}(u,v^{\prime})+Z_{h}(v^{\prime},u)$ for any $v^{\prime}\sim u$ so $v^{\prime}$ is not matched to $u$. Reciprocally, if $u$ is not matched by $\mathbb{M}_{h}$, then
$\max(0,\max_{\begin{subarray}{c}v^{\prime}\sim u\\ v^{\prime}\neq v\end{subarray}}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))=Z_{h}(v^{\prime},u)\geq w(u,v^{\prime})-Z_{h}(u,v^{\prime})$
for any $v^{\prime}\sim u$, so the maximum of $\operatorname*{argmax}_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))$ is not reached by any $v^{\prime}\sim u$ or there are several $v^{\prime}$ in the $\operatorname*{argmax}$, since $\omega$ is non-atomic there is almost surely no ties and the $\max$ is not reached, so $0\geq\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})$ which concludes.
iii) The goal is to show that the law of $Z_{h}$ restricted to a $H-$neighbourhood of $(o_{+},v)$ with $v$ chosen uniformly among the children of $o_{+}$ is the same as the law of $Z_{h}$ restricted to the $H-$neighbourhood of $(o_{-},o_{+})$. Since Recursion (3) is preserved, one only needs to show that the exterior variables of $Z_{h}$ on the $H-$boundary of $(o_{+},v)$ are i.i.d variables of law $\zeta_{h}$. The $H-$neighbourhood of $(o_{+},v)$ is included in the $H+1-$neighbourhood of $(o_{-},o_{+})$ and can be computed from it. One can check by applying Recursion (3) up to twice from the $H+1$-boundary of $(o_{-},o_{+})$, that we effectively recover variables $Z$ on the $H-$boundary of $(o_{+},v)$ that are i.i.d of law $\zeta_{h}$. $(\mathbb{T},Z_{h})$ is thus stationary.
Clearly, the law of the $Z_{h}$ are also symmetric so changing $(o_{-},o_{+})$ into $(o_{+},o_{-})$ does not change their laws, $(\mathbb{T},Z_{h})$ is thus reversible. ∎
From now on, despite resampling $\mathbb{T}$ when applying Proposition 9, we will forgo the notation $\mathbb{T}^{\prime}$ for $\mathbb{T}$ to ease notations.
We conclude this Section with a technical result we will need later. It deals with a property of the probability distributions which are solutions of Equation (4).
Lemma 10.
@@
$h(0)=\hat{\phi}(1-\mathbb{E}_{W\sim\omega}[h_{-}(W)]).$
If $\hat{\phi}(0)>0$ we are done, else, assume $\hat{\phi}(0)=0$ and $h(0)=0$.
Because $\hat{\phi}>0$ on ]0,1]]0,1]] 0 , 1 ], we deduce that $\mathbb{E}_{W\sim\omega}[h_{-}(W)]=1$. This implies in particular that $\inf(\mathrm{supp}(\omega))\geq\sup(\mathrm{supp}(\zeta_{h}))$. Since $\mathrm{supp}(\zeta_{h})\subset\mathbb{R}_{+}$, this implies that $\mathrm{supp}(\omega)\subset\mathbb{R}_{+}$ aswell.
@@
$h(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W-t)\right]\right).$
@@
$h(W^{\prime})=\mathbbm{1}_{W^{\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W-W^{\prime})\right]\right)$
We’ve shown previously that $W^{\prime}\geq 0$ almost surely so we can discard $\mathbbm{1}_{W^{\prime}\geq 0}$. Now take expectation with respect to $\mathcal{W}^{\prime}$, this yields:
$1=\mathbb{E}_{W^{\prime}\sim\omega}\left[\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W-W^{\prime})\right]\right)\right].$
@@
$\mathbb{E}_{W\sim\omega}[h_{-}(W-W^{\prime})]=0.$
Let $W^{\prime},W^{\prime\prime}$ independent variables each of law $\omega$, evaluate equation (9) on $W^{\prime}-W^{\prime\prime}$:
$h(W^{\prime}-W^{\prime\prime})=\mathbbm{1}_{W^{\prime}-W^{\prime\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W+W^{\prime\prime}-W^{\prime})\right]\right).$
Taking expectations:
$0=\mathbb{E}\left[\mathbbm{1}_{W^{\prime}-W^{\prime\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W-(W^{\prime}-W^{\prime\prime}))\right]\right)\right].$
So almost surely, either $W^{\prime}-W^{\prime\prime}<0$ or $\mathbb{E}_{W\sim\omega}[h(W-(W^{\prime}-W^{\prime\prime}))]=1$. Conditionally on $W^{\prime}-W^{\prime\prime}\geq 0$, $\mathbb{E}_{W\sim\omega}[h(W-(W^{\prime}-W^{\prime\prime}))]=1$. But $h_{-}$ is non-decreasing so this statement stays true for $W^{\prime}-W^{\prime\prime}<0$.
So now, taking expectations on $W^{\prime}-W^{\prime\prime}$, we obtained:
$\mathbb{E}(h_{-}(W+W^{\prime\prime}-W^{\prime}))=1.$
We can now then iterate: Let $(W_{i})_{i\in\mathbb{N}^{*}}$ i.i.d variables of law $\omega$, we will show the following statement by induction:
$\forall n\in\mathbb{N}^{*},\mathbb{E}\left[h_{-}\left(\sum_{i=1}^{n}W_{2i-1}-\sum_{i^{\prime}=1}^{n-1}W_{2i}\right)\right]=1.$ (10)
The statement holds for $n=2$, assume it holds for some $n\in\mathbb{N}^{*}$. Evaluate $h$ at $\sum_{i=1}^{n}W_{2i-1}-\sum_{i^{\prime}=1}^{n-1}W_{2i}$, this yields
$h\left(\sum_{i=1}^{n}W_{2i-1}-\sum_{i^{\prime}=1}^{n-1}W_{2i}\right)=\mathbbm{1}_{\sum_{i=1}^{n}W_{2i-1}-\sum_{i^{\prime}=1}^{n-1}W_{2i}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]\right).$
where the expectation is over $W_{2n}$.
Now take expectation over $(W_{i})_{i\leq 2n-1}$, by hypothesis the left-hand side will evaluate to one:
$1=\mathbb{E}\left[\mathbbm{1}_{\sum_{i=1}^{n}W_{2i-1}-\sum_{i^{\prime}=1}^{n-1}W_{2i}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h_{-}\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]\right)\right].$
@@
$\mathbb{E}_{W_{2n}\sim\omega}\left[h_{-}\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]=0.$
Now evaluate $h$ on $\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}$, this yields:
$h$ $\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)$
$\quad=\mathbbm{1}_{\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h_{-}\left(W_{2n+1}-\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right)\right]\right),$
@@
$0=\mathbb{E}\left[\mathbbm{1}_{\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h_{-}\left(W_{2n+1}-\left(\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right)\right]\right)\right].$
@@
$\mathbb{P}\left(\mathbb{E}_{W_{2n+1}\sim\omega}\left[h_{-}\left(\sum_{i^{\prime}=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{2i}\right)\right]=1\middle|\sum_{i^{\prime}=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0\right)=1.$
Again, since $h_{-}$ is non-decreasing, we can discard the conditioning to obtain:
$\mathbb{P}\left(\mathbb{E}_{W_{2n+1}\sim\omega}\left[h_{-}\left(\sum_{i^{\prime}=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{2i}\right)\right]=1\right)=1.$
@@
$\mathbb{E}\left[h_{-}\left(\sum_{i^{\prime}=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{2i}\right)\right]=1.$
As required.
Now Equation (10) implies that:
$\forall n\in\mathbb{N}^{*},\inf(\mathrm{supp}(\omega))-n\left[\sup(\mathrm{supp}(\omega))-\inf(\mathrm{supp}(\omega))\right]\geq\sup(\mathrm{supp}(\zeta_{h})).$
But since $\omega$ is atomless, $\mathrm{supp}(\omega)$ is not reduced to a single point so $\sup(\mathrm{supp}(\omega))-\inf(\mathrm{supp}(\omega))>0$. Taking $n\rightarrow\infty$ then shows that $-\infty\geq\sup(\mathrm{supp}(\zeta_{h}))$, this is clearly impossible as $\zeta_{h}$ is a law on $\mathbb{R}_{+}$.
Now let us show that $\zeta$ is non-atomic outside of zero, this is essentially a consequence of the RDE, for $Z,Z_{i}\sim\zeta$ and $w_{i}\sim\omega$, all mutually independent :
@@ 3.2 Adding self-loops
In the remainder of this section, we will adopt the vertex-rooted point of view.
[…]
This fact may seem quite surprising at first, but when we come back to the finite case, it is really the atomless property of the weights that makes it so that the optimal matching is unique almost surely. In the infinite case, this still holds.
We will first introduce an extension of $\mathbb{T}$ as a graph where optimality of matchings on $\mathbb{T}$ transfers over. The main reason that led us to do this is to look at the vertex-rooted definition of performance:
$\operatorname{perf_{V}}(\mathbb{T},\mathbb{M})=\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\right]$
We would like to use the recursion on $Z_{h}$ to compare $\mathbb{M}_{h}$ to other matchings $\mathbb{M}$. In the previous expression, there are two cases depending on whether $o$ is matched:
  • If $o$ is matched to some $v$, $w(o,v)$ intervenes in the recursion of $Z_{h}(u,o)$ and $Z_{h}(u,v)$ so the information transported by the $Z_{h}$ has "compared" $w(o,v)$ with other quantities, we can then hope to use the recursion to study the optimality of $v$.
  • However, if the sum is empty, the contribution is 00 which has not intervened anywhere, so we would a priori have to study this exception everywhere.
This introduces very cumbersome exceptional cases that become intractable when we start to consider complex events, the goal is thus to extend the weights and make the random variables $Z_{h}$ on self edges $(v,v)$ so that the weights on $(v,v)$ (that is zero for now, but we will see that we will put non zero values on them instead) are transported through the recursion as well. The definition of $Z_{h}(u,u)$ that we introduce shortly after also has a simple interpretation, it is the (signed) gain if one forced u to be matched to one of its neighbours.
Another more technical motivation to do this is to recover the same recursion that has been studied by Aldous in [aldous2000zeta2].
We will define the extension $(\mathbb{T}^{s},Z_{h}^{s},(f_{i}^{s})_{i\in\{0,...,I\}})=((V^{s},E^{s}),o,w^{s},Z_{h}^{s},(f_{i}^{s})_{i\in\{0,....,I\}})$ deterministically on every outcome of $(\mathbb{T},Z_{h})=((E,V),o,w,Z_{h},(f_{i})_{i\in\{0,...,I\}})$, refer to Figure 7 for an illustration. The vertex set and the decorations remain unchanged, $V^{s}=V,f_{i}^{s}=f_{i}$. We will enrich the edge set by adding self-loops $(v,v)$ for every $v\in V$ to construct $E^{s}$. If $e=(u,v)$ with $u\neq v$, we do not change $w$ nor $Z_{h}$: we set $w^{s}(u,v)=w(u,v)$ and $Z_{h}^{s}(u,v)=Z_{h}(u,v)$.
@@
$w^{s}(v,v)=Z_{h}^{s}(v,v).$
The following recursive equation holds for $(\mathbb{T}^{s},Z_{h}^{s})$:
$\forall(u,v)\in\overset{\rightarrow}{E^{s}},Z_{h}^{s}(u,v)=\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}(w^{s}(v,u^{\prime})-Z_{h}^{s}(v,u^{\prime})).$ (11)
Indeed if $u=v$ then it is obvious by definition, now if $u\neq v$, then by the previous equation:
$Z_{h}^{s}(u,v)$ $=Z_{h}(u,v)=\max(0,\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u,v\end{subarray}}(w(v,u^{\prime})-Z_{h}(v,u^{\prime})))$
$=\max(w^{s}(v,v)-Z_{h}^{s}(v,v),\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u,v\end{subarray}}(w(v,u^{\prime})-Z_{h}(v,u^{\prime})))$
$=\max(w^{s}(v,v)-Z_{h}^{s}(v,v),\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u,v\end{subarray}}(w^{s}(v,u^{\prime})-Z_{h}^{s}(v,u^{\prime})))$
$=\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}(w^{s}(v,u^{\prime})-Z_{h}^{s}(v,u^{\prime}))$
Now, since $v$ is unmatched by $\mathbb{M}_{h}$ is equivalent to $\max(0,\max_{v^{\prime}\sim v}w(v,v^{\prime})-Z_{h}(v,v^{\prime}))$ being zero from Proposition 9 ii) and $Z_{h}^{s}(v,v)=\max_{v^{\prime}\sim v}w(v,v^{\prime})-Z_{h}(v,v^{\prime})$. We deduce that $(v,v)\in\mathbb{M}_{h}^{s}$ if and only if $w^{s}(v,v)=Z_{h}^{s}(v,v)\leq 0$. Since the $w$’s are non-atomic, we can replace the previous inequality with a strict one. As a consequence, the implication we defined previously for the criterion of edges being in $\mathbb{M}_{h}$ is now an equivalence:
$\{u,v\}\in\mathbb{M}_{h}^{s}\Leftrightarrow w^{s}(u,v)>Z_{h}^{s}(u,v)+Z_{h}^{s}(v,u)\Leftrightarrow v=\operatorname*{argmax}_{v^{\prime}\sim v}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})).$ (12)
This is the part that would break down had we chosen $\omega$ to be atomic, the last inclusion rule would become fuzzy with multiple ties.

### v2

@@ 3.1 A message passing algorithm
$w(u,v)>Z_{h}(u,v)+Z_{h}(v,u)\Rightarrow\{v\}=\operatorname*{argmax}_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})).$
@@
To show the equivalent vertex rule, we only need to show that $u$ is not matched if and only if $\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\leq 0$.
For the if part, we have that any incident $Z_{h}(v,u)$ for $v\sim u$ evaluates to zero by applying Recursion (3). So for any $v\sim u$, $Z_{h}(v,u)+Z_{h}(u,v)=Z_{h}(u,v)$. Since $\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\leq 0$, we deduce that for any $v^{\prime}\sim u$, $w(u,v^{\prime})\leq Z_{h}(u,v^{\prime})$. But we have just seen that $Z_{h}(v^{\prime},u)=0$, hence $w(u,v^{\prime})\leq Z_{h}(u,v^{\prime})+Z_{h}(v^{\prime},u)$ for any $v^{\prime}\sim u$. So $v^{\prime}$ is not matched to $u$.
Reciprocally, if $u$ is not matched by $\mathbb{M}_{h}$, then for any $v^{\prime}\sim u$
$\max(0,\max_{\begin{subarray}{c}v^{\prime}\sim u\\ v^{\prime}\neq v\end{subarray}}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})))=Z_{h}(v^{\prime},u)\geq w(u,v^{\prime})-Z_{h}(u,v^{\prime}),$
so the maximum of $\operatorname*{argmax}_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))$ is either not reached by any $v^{\prime}\sim u$ or there are several $v^{\prime}$ in the $\operatorname*{argmax}$. Since $\omega$ is non-atomic, there is almost surely no ties and the $\max$ is not reached. So $\max_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime})\leq 0$ which concludes. ∎
From now on, despite resampling $\mathbb{T}$ when applying Proposition 3.2, we will forgo the notation $\mathbb{T}^{\prime}$ for $\mathbb{T}$ to ease notations.
We conclude this section with a technical, but useful, result. It gives a property of the probability distributions that are solutions of Equation (4).
Lemma 3.3.
@@
$h(0)=\hat{\phi}(1-\mathbb{E}_{W\sim\omega}[h(W)]).$
The result if immediate if $\hat{\phi}(0)>0$, hence we assume $\hat{\phi}(0)=0$ and $h(0)=0$. Since $\hat{\phi}>0$ on ]0,1]]0,1]] 0 , 1 ], we deduce that $\mathbb{E}_{W\sim\omega}[h(W)]=1$. This implies in particular that $\inf(\mathrm{supp}(\omega))\geq\sup(\mathrm{supp}(\zeta_{h}))$. Since $\mathrm{supp}(\zeta_{h})\subset\mathbb{R}_{+}$, this implies that $\mathrm{supp}(\omega)\subset\mathbb{R}_{+}$ as well.
@@
$h(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h(W-t)\right]\right).$
@@
$h(W^{\prime})=\mathbbm{1}_{W^{\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h(W-W^{\prime})\right]\right)$
We have shown previously that $W^{\prime}\geq 0$ almost surely so we can discard $\mathbbm{1}_{W^{\prime}\geq 0}$. Now take expectation with respect to $W^{\prime}$, this yields:
$1=\mathbb{E}_{W^{\prime}\sim\omega}\left[\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h(W-W^{\prime})\right]\right)\right].$
@@
$\mathbb{E}_{W\sim\omega}[h(W-W^{\prime})]=0.$
Let $W^{\prime},W^{\prime\prime}$ independent variables of law $\omega$. Evaluate equation (10) on $W^{\prime}-W^{\prime\prime}$:
$h(W^{\prime}-W^{\prime\prime})=\mathbbm{1}_{W^{\prime}-W^{\prime\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h(W+W^{\prime\prime}-W^{\prime})\right]\right).$
Taking the expectation on both sides,
$0=\mathbb{E}\left[\mathbbm{1}_{W^{\prime}-W^{\prime\prime}\geq 0}\hat{\phi}\left(1-\mathbb{E}_{W\sim\omega}\left[h(W-(W^{\prime}-W^{\prime\prime}))\right]\right)\right].$
So almost surely, either $W^{\prime}-W^{\prime\prime}<0$ or $\mathbb{E}_{W\sim\omega}[h(W-(W^{\prime}-W^{\prime\prime}))]=1$. Conditionally on $W^{\prime}-W^{\prime\prime}\geq 0$, $\mathbb{E}_{W\sim\omega}[h(W-(W^{\prime}-W^{\prime\prime}))]=1$. But the function $h$ is non-decreasing so this statement stays true for $W^{\prime}-W^{\prime\prime}<0$.
So now, taking expectations on $W^{\prime}-W^{\prime\prime}$, we obtain:
$\mathbb{E}(h(W+W^{\prime\prime}-W^{\prime}))=1.$
We can now iterate this procedure. Let $(W_{i})_{i\in\mathbb{N}^{\star}}$ be i.i.d variables with law $\omega$, we will show the following statement by induction:
$\forall n\in\mathbb{N}^{\star},\mathbb{E}\left[h\left(\sum_{i=1}^{n}W_{2i-1}-\sum_{i=1}^{n-1}W_{2i}\right)\right]=1.$ (11)
The statement holds for $n=2$. Assume it holds for some $n\in\mathbb{N}^{\star}$. Evaluating $h$ at $\sum_{i=1}^{n}W_{2i-1}-\sum_{i=1}^{n-1}W_{2i}$, we get
$h\left(\sum_{i=1}^{n}W_{2i-1}-\sum_{i=1}^{n-1}W_{2i}\right)=\mathbbm{1}_{\sum_{i=1}^{n}W_{2i-1}-\sum_{i=1}^{n-1}W_{2i}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]\right).$
where the expectation is taken only over $W_{2n}$.
Now take expectation over $(W_{i})_{i\leq 2n-1}$, by the induction hypothesis the left-hand side will evaluate to one:
$1=\mathbb{E}\left[\mathbbm{1}_{\sum_{i=1}^{n}W_{2i-1}-\sum_{i=1}^{n-1}W_{2i}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]\right)\right].$
@@
$\mathbb{E}_{W_{2n}\sim\omega}\left[h\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right]=0.$
Now evaluate $h$ on $\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}$, this yields:
$h$ $\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)$
$\quad=\mathbbm{1}_{\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h\left(W_{2n+1}-\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right)\right]\right),$
@@
$0=\mathbb{E}\left[\mathbbm{1}_{\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0}\hat{\phi}\left(1-\mathbb{E}\left[h\left(W_{2n+1}-\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right)\right]\right)\right].$
@@
$\mathbb{P}\left(\mathbb{E}_{W_{2n+1}\sim\omega}\left[h\left(W_{2n+1}-\left(\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\right)\right)\right]=1\middle|\sum_{i=1}^{n}W_{2i}-\sum_{i=1}^{n}W_{2i-1}\geq 0\right)=1.$
Again, since $h$ is non-decreasing, we can discard the conditioning to obtain:
$\mathbb{P}\left(\mathbb{E}_{W_{2n+1}\sim\omega}\left[h\left(\sum_{i=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{2i}\right)\right]=1\right)=1.$
@@
$\mathbb{E}\left[h\left(\sum_{i=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{2i}\right)\right]=1,$
which completes the induction. Now, introducing a variable $Z_{h}$ of law $\zeta_{h}$ independent from the $(W_{i})_{1\leq i\leq 2n+1}$, Equation (11) can be rewritten as:
$\forall n\in\mathbb{N}^{*},\mathbb{P}\left(Z_{h}\leq\sum_{i=1}^{n+1}W_{2i-1}-\sum_{i=1}^{n}W_{i}\right)=1.$
This implies that:
$\forall n\in\mathbb{N}^{\star},\inf(\mathrm{supp}(\omega))-n\left[\sup(\mathrm{supp}(\omega))-\inf(\mathrm{supp}(\omega))\right]\geq\sup(\mathrm{supp}(\zeta_{h})).$
But since $\omega$ is atomless, $\mathrm{supp}(\omega)$ is not reduced to a single point so $\sup(\mathrm{supp}(\omega))-\inf(\mathrm{supp}(\omega))>0$. Taking $n\rightarrow\infty$ then shows that $-\infty\geq\sup(\mathrm{supp}(\zeta_{h}))$. This is clearly impossible as $\zeta_{h}$ is a law on $\mathbb{R}_{+}$. We then conclude that $h(0)>0$ which exactly means that $\zeta_{h}$ is atomic at zero.
Now let us observe that $\zeta$ is non-atomic outside of zero. This is essentially a consequence of the RDE, for $Z,Z_{i}\sim\zeta$ and $w_{i}\sim\omega$, all mutually independent:
@@ 3.2 Adding self-loops
In the remainder of this section, we will adopt the vertex-rooted point of view. Dealing with partial matchings rather than perfect matchings will be very cumbersome in future proofs. For example, when comparing two matchings, we would need to discuss several cases depending on whether a vertex is matched or not for the two matchings. A simple solution to deal with this is to add self loops, in which case the graphs always have perfect matchings, and any partial matching on the graph without self loops can be augmented into a perfect matching on the graph with self loops. Reciprocally, a perfect matching on a graph with self loops can be restricted to a partial matching on the corresponding graph without self loops.
We need to choose adequate weights for the self loops so that a partial matching on the graph without self loops is optimal among all partial matchings of the original graph if and only if the corresponding perfect matching on the augmented graph with self loops is optimal among all perfect matchings of the augmented graph.
Starting with a system of variables $Z(u,v)$ for $(u,v)\in\overset{\rightarrow}{E}$ that satisfies recursion (3), we want to extend it to variables $Z(v,v)$ for each self loop so that we get a similar recursion as (3) and the decision rule (2) constructs the perfect matching of the augmented graph associated to the matching defined by the original variables. Recall that, for $u\neq v$ and $u\sim v$, the original recursion (3) on the variables $Z$ reads
$Z(u,v)=\max(0,\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u,v\end{subarray}}(w(v,u^{\prime})-Z(v,u^{\prime}))).$
To take into account the self loop $(v,v)$, we want instead
$Z(u,v)=\max(0,\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}(w(v,u^{\prime})-Z(v,u^{\prime}))).$
Choosing
$w(v,v)=Z(v,v)$
ensures that this recursion is still valid and reads
$Z(u,v)=\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}(w(v,u^{\prime})-Z(v,u^{\prime})).$
For $Z(v,v)$ to satisfy the same recursion, we set
$w(v,v)=Z(v,v):=\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq v\end{subarray}}(w(v,u^{\prime})-Z(v,u^{\prime})).$
This definition of $Z(v,v)$ has a simple interpretation in the finite setting: it is the (signed) gain if one forces $v$ to be matched to one of its neighbours.
Formally, we define the extension
$(\mathbb{T}^{s}_{h},Z_{h}^{s},(f_{i}^{s})_{i\in\{0,...,I\}})=((V^{s},E^{s}),o,w^{s}_{h},Z_{h}^{s},(f_{i}^{s})_{i\in\{0,....,I\}})$ (12)
deterministically on every outcome of $(\mathbb{T},Z_{h})=((V,E),o,w,Z_{h},(f_{i})_{i\in\{0,...,I\}})$, refer to Figure 7 for an illustration. The vertex set, the root and the decorations remain unchanged, $V^{s}=V$ $o^{s}=o$ and $f_{i}^{s}$ restricted to $\overset{\rightarrow}{E}$ is $f_{i}$ and +$\infty$+ ∞ otherwise. We will enrich the edge set by adding self-loops $(v,v)$ for every $v\in V$ to construct $E^{s}$. If $e=(u,v)$ with $u\neq v$, we do not change $w$ nor $Z_{h}$: we set $w^{s}_{h}(u,v)=w(u,v)$ and $Z_{h}^{s}(u,v)=Z_{h}(u,v)$.
@@
$w^{s}_{h}(v,v)$ $=Z_{h}^{s}(v,v).$
The following recursive equation holds for $(\mathbb{T}^{s}_{h},Z_{h}^{s})$:
$\forall(u,v)\in\overset{\rightarrow}{E^{s}},\quad Z_{h}^{s}(u,v)=\max_{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}(w^{s}_{h}(v,u^{\prime})-Z_{h}^{s}(v,u^{\prime})).$ (13)
Remark.
This is the same recursion that has been studied by Aldous [aldous2000zeta2] in the setting of the complete graph and the Poisson Weighted Infinite Tree.
Now, let us look at the analogous decision to the decision rule (2):
$\{u,v\}\in\mathbb{M}_{h}^{s}\Leftrightarrow w^{s}_{h}(u,v)>Z_{h}^{s}(u,v)+Z_{h}^{s}(v,u).$
and show that $\mathbb{M}_{h}^{s}$ is the perfect matching of $\mathbb{T}^{s}_{h}$ corresponding to $\mathbb{M}_{h}$.
If $(u,v)$ is not a self loop, the rule for $\mathbb{M}_{h}^{s}$ is the same as for $\mathbb{M}_{h}$ since we have not modified the variables appearing in the equation. Therefore, the only extra edges of $\mathbb{M}_{h}^{s}$ are self loops.
Now we need to show that a self loop $(v,v)\in\mathbb{M}_{h}^{s}$ iff the vertex $v$ is not matched in $\mathbb{M}_{h}$. By Proposition 3.2 ii), $v$ is unmatched by $\mathbb{M}_{h}$ if and only if $\max(0,\max_{v^{\prime}\sim v}w(v,v^{\prime})-Z_{h}(v,v^{\prime}))=0$. Since we have defined $Z_{h}^{s}(v,v)$ as $\max_{v^{\prime}\sim v}w(v,v^{\prime})-Z_{h}(v,v^{\prime})$, we deduce that $v$ is unmatched by $\mathbb{M}_{h}$ if and only if $w^{s}_{h}(v,v)=Z_{h}^{s}(v,v)\leq 0$. Since the $w$’s are non-atomic, we can replace the previous inequality with a strict one. On the other hand, the rule for self loops reads
$\{v,v\}\in\mathbb{M}_{h}^{s}\Leftrightarrow w^{s}_{h}(v,v)>Z_{h}^{s}(v,v)+Z_{h}^{s}(v,v)=2\,w^{s}_{h}(v,v),$
which is equivalent to $w^{s}_{h}(v,v)<0$ and to $v$ not being matched in $\mathbb{M}_{h}$.
Furthermore we have also found the equivalent vertex-rule for $\mathbb{M}_{h}^{s}$:
$\{u,v\}\in\mathbb{M}_{h}^{s}\Leftrightarrow w^{s}_{h}(u,v)>Z_{h}^{s}(u,v)+Z_{h}^{s}(v,u)\Leftrightarrow v=\operatorname*{argmax}_{v^{\prime}\sim v}(w^{s}_{h}(u,v^{\prime})-Z_{h}^{s}(u,v^{\prime})).$ (14)